amik:

A 33 gram sample of a substance that's a by-product of fireworks has a k-value of 0.1081. N = N0e -kt N0 = initial mass (at time t = 0) N = mass at time t k = a positive constant that depends on the substance itself and on the units used to measure time t = time, in days Find the substance's half-life, in days. Round your answer to the nearest tenth

Timmyspu:

What is the first thing we have to do here?

Timmyspu:

Also welcome to question cove.

SemiDefinite:

to find the 1/2 life: \(\dfrac{N}{N_o} = \dfrac{1}{2}\)